# From strong to weak correlations in breathing-mode kagome van der Waals materials: Nb3(F,Cl,Br,I)8 as a robust and versatile platform for many-body engineering

### v1

@@ II.5.1 Half-Filling ($N_{e}=2$)
We note that the correlation induced narrowing of the band-width in $\Gamma$-$A$ direction within our cluster DMFT calculations in comparison to our DFT calculations is on the order of $30$ to 505050\,50meV and thus likely below current experimental accuracies. Also, the DFT band-width might be underestimated as a result of the applied PBE functional. Further comparison with the available ARPES results show that the lower $t^{2}_{2g}$ bands, which we neglect within our cluster DMFT calculations, are still separated from the upmost bands by about $0.5$ to 111\,1eV (increasing from I to Br), which is in agreement with the trends expected from comparison to DFT calculations, ruling out charge-transfer insulating states.
Based on these observations it is delicate to unambiguously differentiate between conventional band, strongly correlated, and Mott insulators based on ARPES data. Recently Date et al. [24] suggested to analyze details of the upmost (lower Hubbard) band and its $k_{z}$ dispersion to differentiate between the Mott and band insulating scenario in Nb3Br8. Based on our cluster DMFT calculations, we suggest to furthermore look for the existence of the secondary LHB. To this end, we analyzed our EDCs in detail, see Fig. 7 (e-h) as well as Appendix .8. We find for all photon energies that a good fit for Nb3Br8 requires taking a third band at lower energies into account. We note that this feature might also be present in the EDC curve fits by Date et al. [24]. In contrast, for Nb3I8 we do not consistently observe this third feature (as discussed in more detail in Appendix .8) and it is absent in the 58 eV data, Fig. 7 (d). This is in agreement with our cluster DMFT calculations that predict a significant spectral weight reduction of the secondary LHB in Nb3I8 in contrast to Nb3Br8 and is further stressed in the statistical analysis we present in Appendix .8.
II.5.2 Three-Quarter Filling ($N_{e}=3$)
Figure 8: Theoretical Spectral Functions at $N_{e}=3$. (a-d) Interacting spectral functions of doped compounds ($N_{e}=3$) calculated using the Hubbard-I approximation. (e) Spectral functions of the local Green’s functions.
To uncover the degree of correlations, we further suggest studying the doping dependence of the different compounds, which could serve as a direct and unambiguous tool to differentiate between strongly and weakly correlated insulators. To this end, we start with doping the system by exactly one electron (or hole) per dimerized inter-layer trimer. Experimentally, this could be achieved by replacing one halide with one chalcogen atom per dimerized unit cell, e.g., Nb3X7.5Te0.5 (note that Refs. [53] and [54] find Nb3Cl7Te and Nb3I7Te still displaying the breathing mode kagome lattice, respectively). In this case we still deal with integer filling of the two flat bands, but with $N_{e}=3$ instead of half filling with $N_{e}=2$, such that we can still use the Hubbard-I approximation. The resulting doped spectral functions are shown in Fig. 8. From a pure band-structure perspective, we would expect that doping the systems with one electron (hole), would just half-occupy the non-interacting anti-bonding (boding) band, making all systems good metals. Instead, we see that all systems have a gap of around 0.50.50.5\,0.5eV, which we understand as clear evidence for strong correlation effects developing in all systems, including Nb3I8. In detail, by doping with only one electron (hole), Nb3F8* and Nb3Cl8 are now strongly correlated charge transfer insulators (see overlap between the new uncorrelated bonding bands and the lower Hubbard band of the anti-bonding band), while Nb3Br8 and Nb3I8 are now Mott insulators. In all cases the gaps are opened by strong retardation effects of the self-energy as we explain in more detail in Appendix .5. Taking the approximate particle-hole symmetry of the flat non-interacting bonding/anti-bonding bands into account, we expect that doping in the other direction, to $N_{e}=1$, yields similar correlated electronic structures. We further note that for $N_{e}=3$, the non-local (density-density) Coulomb interaction is playing a quantitatively important role by significantly increasing the gap. If we neglect this contribution, the strongly correlated gaps are much smaller and even close in Nb3F8* (see Appendix .3).
Figure 9: Doping Dependence. Behavior under doping of (a) the quasiparticle weight of the bonding (solid) and anti-bonding (dashed) components and the spectral functions of (b) Nb3Cl8 and (c) Nb3I8. Squares in (a) indicate doping levels, which are analysed in (b) and (c).
II.6 Continuous Doping ($N_{e}=2+\delta$)
To analyze the details of the changes from $N_{e}=2$ to $N_{e}=3$, we performed additional calculations for the Cl, Br, and I compounds at various doping levels in between denoted by $N_{e}=2+\delta$. In these cases Hubbard-I is not a good approximation any more, such that we are restricted to imaginary time/frequency cluster DMFT calculations (see Methods for details). The resulting imaginary frequency self-energies are discussed in Appendix .4. In the bonding orbital channel we find for small dopings $\delta$ stronger dynamic correlation effects in Nb3Cl8 and Nb3Br8 compared to Nb3I8, which become rather weak for all materials toward $N_{e}=3$. The anti-bonding self-energy behaves oppositely for all systems. Upon increasing the doping from $N_{e}=2$ to $N_{e}=3$, dynamic correlations generally increase. This is also reflected in the orbital-dependent quasiparticle weights, which we extracted from the imaginary frequency self-energies and show in Fig. 9 (a). The bonding quasiparticle weights of Nb3Cl8 and Nb3Br8 start close to zero in the vicinity of $N_{e}=2$, indicative of strong correlation effects, and steadily increase with doping, while for Nb3I8 it is larger than $0.5$ at all fillings, indicative of moderate correlation effects. On the other hand, the anti-bonding orbital quasiparticle weight drastically decreases towards $N_{e}=3$ for all three compounds. This can be understood from the band perspective, where $N_{e}=3$ corresponds to a half-filled anti-bonding band. The large Coulomb interaction compared to the kinetic energy thus renders the $N_{e}=3$ systems more or less conventional single-band Mott insulators, but with a lower Hubbard band that is close to the fully-occupied bonding band.
To gain further spectroscopic insights into the continuous doping responses of the systems, we continued our imaginary frequency data to the real axis. To this end, we apply the recently proposed minimal pole representation method  [55, 56] to analytically continue the matrix-valued local self-energies and calculate from this the interacting local spectra for Nb3Cl8 and Nb3I8. We chose the latter two to highlight the differences between the doping dependence of a strongly and a weakly correlated system. The results are shown in Fig. 9 (b) and (c). For Nb3Cl8 starting with a small doping of $\delta=0.05$, we see that the main features shift and, importantly, a small resonance at $\omega=0\,$eV arises, which grows in intensity towards $\delta=0.3$, as indicated by the red arrows in Fig. 9 (b). This is reminiscent of a Kondo resonance, which is a hall-mark of doped strongly correlated insulators. Simultaneously, we see that at higher energies new resonances appear, which form at large dopings ($\delta=0.87$) a new upper Hubbard band. In fact, we can understand the $\delta=0.87$ data as a slightly hole-doped variant of the $N_{e}=3$ situations, which already shows the single-orbital lower and upper Hubbard bands, but here, due to finite hole doping, accompanied by a Kondo resonance, as indicated by the green arrow in Fig. 9 (b).
In contrast, the Nb3I8 spectra more or less just shift in energy until around $\delta=0.35$, without introducing any new features and especially without showing any evidence of a Kondo resonance at zero energy. This is the doping dependence we expect from conventional band insulators. However, for larger dopings, we start to see new features, which are in line with our interpretation of Nb3I8 as a single-band Mott insulator at $N_{e}=3$. In detail, at $\delta=0.9$ we should understand Nb3I8 as a hole-doped single-band Mott insulator.
II.7 Quantum Magnetic Properties as the Origin of the Lattice Phase Transition
Magnetism is a ubiquitous concept in correlated materials and has correspondingly received considerable attention in Nb3X8 as well. For the Nb3Cl8 monolayer, our previous work [23] and recent spin-spiral DFT [26] calculations both show a magnetically ordered state with 120 angles between the trimer moments. For the bulk LT phase, on the other hand, strong out-of-plane dimerization leads to singlet-formation [29, 35, 22, 21], which is common in bilayer Hubbard models at large $U$ and large $t_{\perp}/t_{\parallel}$ [57, 58, 52].
Although a detailed account of the quantum magnetism in Nb3X8 is beyond the scope of this work, it is insightful to study the connection between the magnetic, electronic and lattice structure that arise from our downfolded models. Given the dominance of the Coulomb interaction over the hopping matrix elements [e.g., for Nb3Br8 $U^{\text{eff}}$ is approximately five (hundred) times larger than the largest out-of-plane (in-plane), see Fig. 5], it is appropriate to describe the magnetic properties of Nb3X8 using a Heisenberg model that incorporates trimer-localized magnetic moments. Accordingly, we approximate the exchange interactions as $J_{ij}=-2(t_{ij})^{2}/U^{\text{eff}}$, where $U^{\text{eff}}$ represents the effective Coulomb interaction, which includes the renormalization of the energy cost due to non-local Coulomb interactions [23]. For more details, see Section V.2. This formulation indicates that the exchange interaction is strongest when the hopping matrix element between two sites is large, leading to antiferromagnetic (AFM) exchange interactions ($J_{ij}<0$). This behavior is characteristic of materials with large $U$, such as all Nb halides, and explains the differences to DFT estimates that do not account for $U$ [59, 26].
Table 2: Molecular orbital effective exchange parameters for bulk Nb3X8 in the LT structure in meV. The labels (s) and (w) denote strongly and weakly bound neighbors in the out-of-plane ($\perp$) directions, respectively, while $n=1,2,3$ corresponds to first, second, and third nearest neighbors in the in-plane ($\parallel$) direction. Note that two values of $J_{\parallel}^{(n)}$ are provided for Nb3Cl8 due to the slight size differences of trimers in adjacent layers, which arise from the R3333 symmetry.
[…]
$J^{(\text{s})}_{\perp}$0.20 -31.92 -62.75 -159.3
$J^{(\text{w})}_{\perp}$0.09 -0.02 -0.93 -2.11
$J^{(1)}_{\parallel}$0.72 -0.45/-1.37 -0.13 -0.41
$J^{(2)}_{\parallel}$0.01 -0.05/-0.06 -0.08 -0.17
$J^{(3)}_{\parallel}$ 0.00 -0.03/-0.02 -0.11 -0.36
The resulting exchange interaction parameters are listed in Table 2. For the nearest-neighbor in-plane interactions $J^{(1)}_{\parallel}$, we do not find a clear trend from Nb3F8* to Nb3I8. For the second- and third-nearest-neighbor in-plane interactions $J^{(2,3)}_{\parallel}$ we, however, see a clear increasing trend from F to I. We further note that $J^{(3)}_{\parallel}$ is (vanishingly) small in Nb3F8* and Nb3Cl8, yet exceeds $J^{(2)}_{\parallel}$ in Nb3Br8 and Nb3I8.
Taking the bulk matrix elements as proxies for their values in the monolayer limits, we expect that all monolayers exhibit purely antiferromagnetic exchange interactions. This leads to a 120 Néel order in Nb3F8* and Nb3Cl8, while Nb3Br8 and Nb3I8 display spin spirals with non-commensurate wave vectors, attributed to the strong exchange interactions between their third nearest neighbors as discussed in more detail in Appendix .7.
@@ II.7 Quantum Magnetic Properties as the Origin of the Lattice Phase Transition
In bulk, the main difference between the LT and HT phases lies in the connectivity of the layers, as illustrated in Fig. 1 (b). In the HT phase, every trimer is connected to three trimers in the layer above and three in the layer below. Because the trimer moments within a layer are coupled antiferromagnetically, with either 120 or incommensurate helical order, this leads to a high degree of magnetic frustration in the HT structure: the trimer-moment cannot align consistently with all three moments in the neighboring layers. On the other hand, in the LT phase, the dimerization in the vertical direction lifts this frustration, every trimer moment has a single trimer moment to which it is coupled most strongly, and a singlet forms. The magnetic frustration in the HT phase is strongest in Nb3I8 and gets weaker towards Nb3Br8 and Nb3Cl8, as visible from the decreasing $J^{(1)}_{\perp}$ listed in Table 2. This trend clearly correlates with the decreasing critical transition temperatures. Based on this observation, we can understand the inter-layer magnetic frustration in the HT phases as a main source for the structural transition.
In Fig. 10, we further show the magnetic susceptibility of both LT and HT phases of Nb3Cl8, obtained from classical Monte Carlo simulations as a function of temperature and magnetic field. For computational details, please refer to Section V.2. The susceptibility in the HT phase conforms to Curie-Weiss, while the structural phase transition to the LT phase has a concomitant collapse of the magnetic susceptibility. This is consistent with available experimental data [30, 29, 35, 31, 40, 25]. Very low temperatures show an upturn of the magnetic susceptibility, but we understand this as the breakdown of the classical simulation since the singlet formation can only be corrected by a full quantum mechanical treatment. Experimentally observed increases of the susceptibility at low temperatures should thus be ascribed to impurities and disorder [29, 35].
@@
Figure 10: Magnetic Susceptibility of Nb3Cl8. Classical Monte Carlo simulations of the magnetic susceptibility for both HT and LT structures of Nb3Cl8 as a function of temperature $T$ and magnetic field $B$. The dashed line marks the transition temperature at $T=90\,\text{K}$.
@@ II.8 Symmetry Breaking in Few-Layer Stacks and Connections to the Field Free Josephson Diode Effect
Figure 11: Symmetries of Finite Stacks of Nb3X8. Inversion (IS) and time-reversal (TRS) symmetry in monolayer (ML), bilayer (BL), trilayer (TL), and quadlayer (QL) of Nb3X8, taking into account various surface (”edge”) terminations. The solid dark red line represents strong interlayer coupling between the nearest trimers, while the dashed line indicates weak coupling. Crosses (in brackets) indicate (possible but not necessary) symmetry-breaking. Red checkmarks denote the preservation of symmetry.
@@
The IS breaking can be straightforwardly discussed: Nb3X8 stacks with an even number of layers are inversion symmetric, while stacks with an odd number of layers break inversion symmetry. In particular, inversion symmetry is broken in the monolayer. The TRS breaking requires more attention, especially with respect to the impact of the dimerization (alternating hybridization) of bilayers, which depends on the number of involved layers. Specifically, there can be top and/or bottom surface layers that only weakly hybridize with their neighbor and can therefore approximately behave like freestanding monolayers. A trilayer, for example, consists of a fully dimerized bilayer accompanied by a nearly-free monolayer, as depicted in Fig. 11. Taking the strong correlation effects into account, these surface monolayers realize (approximately) conventional single-band Mott insulators 222 In the monolayer limit, all four compounds host a single half-filled flat band around the Fermi level and, as a result of reduced screening, local Coulomb interaction matrix elements are larger than in the bulk (e.g., $U=1.9\,$eV for monolayer Nb3Cl8 [23] versus 1.51.51.5\,1.5eV in bulk according to Table 5). For a surface layer in a finite stack, the expectation is that $U$ is between the latter values, since there is screening from one side. , which are only weakly coupled to the neighboring dimerized bilayer, as we show in more detail in Appendix .6. In contrast, mean field DFT calculations, wrongly predict the trilayer to behave like a metallic monolayer in contact with a band-insulating bilayer, cf. Appendix .6 and Ref. [28]. Thus, in the context of symmetry breakings and the Josephson diode effect, correlation properties must be considered consistently throughout the whole structure with a special focus on unpaired surface monolayers.
@@
However, whether a magnetic spin-spiral state and its potential TRS breaking also has a global effect on the electronic properties might depend on the electron coherence length [62]. If the latter is smaller than the spin-spiral wavelength, there might be a global effect. For example, the 120 Néel state of monolayer Nb3Cl8 seems to leave the TRS after backfolding to the primitive unitcell intact [26]. We find the spin-spiral wavelengths to be ranging from 101010\,10Å  to 18.518.518.5\,18.5Å  in Nb3Cl8 and Nb3I8, respectively (see Appendix .7), which is on the same order as the electronic coherence length of other Mott insulators [63, 64, 65]. Based on this, we assume a possible global effect of the TRS breaking in the surface monolayers of Nb3Br8 and Nb3I8, but possibly not in Nb3F8* and Nb3Cl8. Additionally, we note that the interfaces to the superconducting leads in NbSe2/Nb3Br8/NbSe2, as well as magnetic impurities, might further contribute to the TRS breaking.
@@ III Conclusion and Discussion
Our analysis shows that the low-energetic electronic structure as well as the magnetic properties of the whole class of Nb3X8 are dominated by the formation of flat bands as a result of the breathing-mode kagome distortion. Since the resulting low-energetic bands are well separated from those formed by all other Nb and halides orbitals, low energy generalized Hubbard models can be straightforwardly obtained by ab intio downfolding using the constrained random phase approximation in combination with Wannier orbitals representing molecular orbitals centered at the trimerized Nb clusters.
The resulting models allow us to quantify the strengths of in- and out-of-plane hopping as well as local and non-local Coulomb matrix elements. The whole material family is in the regime of strong local interaction ($U_{0}/t_{\parallel}>80$). However, the ratios between these kinetic and local interaction terms suggest that from the putative Nb3F8* to Nb3I8 the degree of Coulomb correlations gradually decrease. Based on a strong coupling expansion we can further derive effective magnetic models, which show that the high-temperature crystal structures host significant inter-layer magnetic frustrations, which are lifted by singlet formation in the low-temperature structures. As these frustrations are the strongest in Nb3I8 and gradually diminish in Nb3Br8 and Nb3Cl8, the accompanying critical temperatures of the crystallographic phase transition also gradually decrease from Nb3I8 to Nb3Cl8.
As a result, the low-temperature bulk materials are best understood as dimerized bilayers, which only couple weakly among each other through vdW interactions. The magnetic ground states are thus most likely singlets formed between inter-layer neighboring spins localized on the strongly hybridized molecular orbitals. On the electronic side, our combined cRPA and cluster DMFT calculations show that the halide atoms control the ratios between the local Coulomb interactions and the in- and out-of-plane hopping terms and thereby drive the bulk materials from a conventional two-band Mott insulator (Nb3F8*) to strongly correlated insulators (Nb3Cl8 and Nb3Br8) and finally to a weakly correlated band insulator (Nb3I8). Thus, already in the undoped case the class of Nb3X8 host several correlated ground states of different characters.
@@
Thus, the combination of in-plane trimerization and out-of-plane dimerization renders Nb3X8 an extremely robust, versatile, but also simple platform to study and engineer many-body phenomena. In combination with the various and tunable degree of correlations, Nb3X8 can thus be seen as a promising drosophila melanogaster of layered correlated (Mott) insulators, which offers the exciting possibility to precisely study numerous correlation phenomena both theoretically and experimentally. The models which we derived here yield an optimal theoretical framework for this purpose, and the complete model specifications for the entire Nb3X8 family are made freely available via a public repository, see Ref. TBA.

### v2

@@ II.7.2 Out-of-plane Magnetic Frustrations as the Origin of the Lattice Phase Transition in Bulk Nb3X8
$J^{(\text{s})}_{\perp}$ -0.03 -35.21 -67.98 -180.21 -0.55 -0.99
$J^{(\text{w})}_{\perp}$ -0.05 -0.49 -0.98 -2.29 -0.48 -0.96
$J^{(1)}_{\parallel}$ -1.32 -0.83 -0.09 -0.26 -1.18 -0.30
$J^{(2)}_{\parallel}$ -0.00 -0.05 -0.07 -0.16 -0.05 -0.07
$J^{(3)}_{\parallel}$ 0.00 -0.04 -0.10 -0.33 -0.04 -0.10
$E_{\text{ex}}$ -3.99 -37.61 -70.38 -185.98 -3.62 -3.94
We proceed with the discussion the magnetic properties in both low- and high-temperature bulk structures as resulting from the magnetic interaction parameters given in Tab. 3.
@@
In bulk, The main difference between the LT and HT phases lies in the connectivity of the layers, as illustrated in Fig. 2 (b). In the HT phase, every trimer is magnetically connected to three trimers in the layer above and three in the layer below by (rather small) $J_{\perp}^{(s/w)}$. As a result of the anti-ferromagnetic inter-layer coupling, this leads to magnetic frustration in the HT structure: the trimer-moment cannot align consistently with all three moments in the upper and lower neighboring layers. On the other hand, in the LT phase, the dimerization in the vertical direction lifts this frustration and every trimer moment has another trimer moment to which it is coupled strongly by $J_{\perp}^{(s)}$, supporting singlet formation. Overall, this lowers the total energy of the system.
@@
Figure 12: Magnetic Susceptibility of Nb3Cl8 and Nb3Br8. Classical Monte Carlo simulations of the magnetic susceptibility for both HT and LT structures as function of temperature $T$ at magnetic field B=5B=5\,T. The dotted vertical lines mark the transition temperatures at $T=92$ and 387387\,K.
This is reflected in the total (magnetic) energies of Nb3Cl8 and Nb3Br8, which are reduced by about 3434\,meV and 6666\,meV in the LT phases, respectively (see Tab. 3). This trend clearly correlates with the critical transition temperatures of Nb3Cl8 and Nb3Br8 [38] and offers a simple explanation of why Nb3I8, that has the largest exchange interaction energy among other LT Nb3X8 compounds, has so far been only observed in the LT phase. These observations made possible by our material-specific matrix elements are in line with the suggestion by Sheckleton et al. [31], that the inter-layer magnetic frustration in the HT phase combined with energy gain due to singlet formation in LT phase drives the structural transition.
We note that in the case of LT Nb3F8*, singlet formation is not energetically favorable. This is due to $J^{\text{(s)}}_{\perp}<J^{\text{(w)}}_{\perp}<J^{\text{(1)}}_{\parallel}$, which indicates that the primary contribution to the ground state arises from the in-plane $120^{\circ}$-AFM ordering.
This interpretation also aligns with the local magnetic susceptibilities, which we compute using classical Monte Carlo simulations (for computational details see Section V.4) for Nb3Cl8 and Nb3Br8 in their LT and HT structures, solving Eq. 2 and considering a finite magnetic field. Indeed, as shown in Fig. 12, the computed magnetic susceptibilities for the HT structures of both Nb3Cl8 and Nb3Br8 match and follow a Curie-Weiss behavior. A concomitant collapse of the magnetic susceptibility accompanies the structural phase transition to the LT phase for both compounds. The reduction in magnetic susceptibility of LT Nb3Br8 compared to LT Nb3Cl8 can be attributed to a larger $J_{\perp}^{\text{(s)}}$, indicating that LT Nb3Br8 is less responsive to magnetic fields. This seems to be perfectly consistent with available experimental data [32, 31, 38, 33, 43, 27]. However, we note that for fully compensated singlets, the susceptibilities in the LT phases should be completely suppressed, while our calculations show small but finite responses accompanied by small tails at low temperatures. This aspect thus necessitates a full quantum mechanical treatment in the future.
@@ II.8 Symmetry Breaking in Few-Layer Stacks and Connections to the Field Free Josephson Diode Effect
Figure 13: Symmetries of Finite Stacks of Nb3X8. Inversion and time-reversal symmetry in monolayer, bilayer, trilayer, and quadlayer of Nb3X8, taking into account various surface (”edge”) terminations. The solid dark red line represents strong interlayer coupling between the nearest trimers, while the dashed line indicates weak coupling. Crosses (in brackets) indicate (possible but not necessary) symmetry-breaking. Red checkmarks denote the preservation of symmetry.
@@
The IS breaking can be straightforwardly discussed: Nb3X8 stacks with an even number of layers are inversion symmetric, while stacks with an odd number of layers break inversion symmetry. In particular, inversion symmetry is broken in the monolayer. The TRS breaking requires more attention, especially with respect to the impact of the dimerization (alternating hybridization) of bilayers, which depends on the number of involved layers. Specifically, there can be top and/or bottom surface layers that only weakly hybridize with their neighbor and can therefore approximately behave like freestanding monolayers. A trilayer, for example, consists of a fully dimerized bilayer accompanied by a nearly-free monolayer, as depicted in Fig. 13. Taking the strong correlation effects into account, these surface monolayers realize (approximately) conventional single-band Mott insulators 222 In the monolayer limit, all four compounds host a single half-filled flat band around the Fermi level and, as a result of reduced screening, local Coulomb interaction matrix elements are larger than in the bulk (e.g., U=1.9U=1.9\,eV for monolayer Nb3Cl8 [24] versus 1.51.5\,eV in bulk according to Tab. LABEL:tab:parameters). For a surface layer in a finite stack, the expectation is that $U$ is between the latter values, since there is screening from one side. , which are only weakly coupled to the neighboring dimerized bilayer, as we show in more detail in Appendix .9. In contrast, mean field DFT calculations, wrongly predict the trilayer to behave like a metallic monolayer in contact with a band-insulating bilayer, cf. Appendix .9 and Ref. [30]. Thus, in the context of symmetry breakings and the Josephson diode effect, correlation properties must be considered consistently throughout the whole structure with a special focus on unpaired surface monolayers.
@@
However, whether a magnetic spin-spiral state and its potential TRS breaking also has a global effect on the electronic properties might depend on the electron coherence length [73]. If the latter is smaller than the spin-spiral wavelength, there might be a global effect. For example, the 120 Néel state of monolayer Nb3Cl8 seems to leave the TRS after backfolding to the primitive unitcell intact [28]. We find the spin-spiral wavelengths to be ranging from 1010\,Å  to 18.518.5\,Å  in Nb3Cl8 and Nb3I8, respectively (see Appendix .10), which is on the same order as the electronic coherence length of other Mott insulators [74, 75, 76]. Based on this, we assume a possible global effect of the TRS breaking in the surface monolayers of Nb3Br8 and Nb3I8, but possibly not in Nb3F8* and Nb3Cl8. Additionally, we note that the interfaces to the superconducting leads in NbSe2/Nb3Br8/NbSe2, as well as magnetic impurities, might further contribute to the TRS breaking.
@@ III Conclusion and Discussion
Our analysis shows that the low-energetic electronic structure as well as the magnetic properties of the whole class of Nb3X8 are dominated by the formation of flat bands as a result of the breathing-mode kagome distortion. Since the resulting low-energetic bands are well separated from those formed by all other Nb and halide orbitals, low energy generalized Hubbard models can be reliably obtained by ab intio downfolding using the constrained random phase approximation in combination with Wannier orbitals representing molecular orbitals centered at the trimerized Nb clusters.
The resulting models properly render the local chemistry of the effective molecular orbitals as well as the global screening properties. This allow us to quantify the strengths of in- and out-of-plane hopping as well as local and non-local Coulomb matrix elements in a material-specific manner. Our ab initio model database further shows that the whole material family is in the regime of strong local interaction ($U_{0}/t_{\parallel}>60$). Since the ratios between these kinetic and local interaction terms vary from monolayer to bulk structures and from Nb3F8* to Nb3I8 as a result of modifications to the global screening, as well as from varying local confinements of the molecular orbitals, the degree of Coulomb correlations gradually decrease from Nb3F8* to Nb3I8 and from ML to bulk structures. Based on a strong coupling expansion we can further derive effective magnetic models, which show that the high-temperature crystal structures host inter-layer magnetic frustrations, which are lifted by singlet formation in the low-temperature structures. As these frustrations are the strongest in Nb3I8 and gradually diminish in Nb3Br8 and Nb3Cl8 accompanied by an decreasing energy gain from singlet formation, the accompanying critical temperatures of the crystallographic phase transition also gradually decrease from Nb3I8 to Nb3Cl8.
As a result, the low-temperature bulk materials are best understood as dimerized bilayers, which only couple weakly among each other through vdW interactions. The magnetic ground states are thus most likely singlets formed between inter-layer neighboring spins localized on the strongly hybridized molecular orbitals. On the electronic side, our combined cRPA and cluster DMFT calculations show that the halide atoms control the ratios between the local Coulomb interactions and the in- and out-of-plane hopping terms and thereby drive the low-temperature bulk materials from a conventional two-band Mott insulator (Nb3F8*) to strongly correlated insulators (Nb3Cl8 and Nb3Br8) and finally to a weakly correlated band insulator (Nb3I8). Moreover, since the screening is significantly reduced in the bilayer structures, we find that even the Nb3I8 bilayer becomes a strongly correlated insulator. In turn, due to the missing out-of-plane dimerization, all monolayer, as well as high-temperature structure variants of Nb3X8, are single- or two-band Mott insulators, respectively. Thus, already in the undoped case the class of Nb3X8 host several correlated ground states of different characters. To differentiate between these variants of strongly and weakly correlated insulators, we showed how the charge-neutral and temperature dependent excitations spectra as well as the charged excitation spectra of the local (interlayer molecular) dimer differ and how this is reflected in experimentally accessible dispersive bonding and nearly non-dispersive anti-bonding lower Hubbard bands.
@@
Thus, the combination of in-plane trimerization and out-of-plane dimerization renders Nb3X8 an extremely robust, versatile, but also simple platform to study and engineer many-body phenomena. In combination with the various and tunable degree of correlations, Nb3X8 can thus be seen as a promising drosophila melanogaster of layered correlated (Mott) insulators, which offers the exciting possibility to precisely study numerous correlation phenomena both theoretically and experimentally. The models which we derived here yield an optimal theoretical framework for this purpose, and the complete model specifications for the entire Nb3X8 family are made freely available via a public repository, see Ref. [86].